OpenStudy (anonymous):

A square pyramid of height 6 inches has a volume of 405 cubic inches. If the volume of a similar square pyramid is 3240 cubic inches, what is its height? 6 inches 12 inches 14 inches 24 inches Could you explain how to solve this? When I tried working this out I got 48.

Directrix (directrix):

@soulflower Apply this theorem: If two solids are similar, the cube of the scale factor of the two solids is equal to the ratio of the volumes. That gives the equation: 405/3240 = (6/h) ^3 where h is the height of the larger similar pyramid. Solve the equation for h. The answer is not 48.

Directrix (directrix):

@soulflower What did you get for h?

OpenStudy (anonymous):

I multiplied 6 and 3240 then divided that by 405 and still got 48. I'm not sure what I'm doing wrong.

Directrix (directrix):

405/3240 = (6/h) ^3 405/3240 simplifies to 1/8 because 405*8 = 3240 So, the equation becomes: 1/8 = (6/h) ^3 (1/2)^3 = (6/h)^3 Take cube roots of both sides and see what you get. @soulflower

Directrix (directrix):

1/2 = 6/h h = ? @soulflower

OpenStudy (anonymous):

12 (: Thank you. I see where I went wrong.

Directrix (directrix):

12 is what I got, too.